# Staking out the Proton Drip-Line of Thulium at the N=82 Shell Closure

### v1

@@ III Discussion
The Er isotopic chain was measured alongside Tm during experiment #2 from 154Er to 149Er.
@@
During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found which significantly affects the one-proton separation energy ($S_{p}$) of the 150Tm ground-state. Together, the newly-measured 150Tm mass and the AME2020 mass of 149gEr results in a $S_{p}$ that is consistent with zero ($1\sigma$), but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm.
@@
As with 151Yb for which the ISAC target was seen to preferentially deliver the spin 11/2 isomer to the ground-state, 149Er, having two fewer protons similarly appeared to prefer the excited state. We therefore fit two hyperEMG functions to the 149Er peak observed at TITAN to account for both the dominant isomeric state ($t_{1/2}=8.9\pm$0.2 s), as well as the ground-state ($t_{1/2}=4\pm$2 s) which shows up as an excess within the low-mass tail.
@@
The ground-state mass of 149Er can be determined by two different methods using the present data. Firstly, it can be established from the measured isomer mass and known excitation energy as was previously done in an experiment using Schottky mass spectrometry [18], which reported an uncertainty of $\pm$28 keV. A second method is to directly fit the ground-state within the tail of the isomer’s peak by first fitting 149Dy to a hyperEMG function to establish the peak shape.
@@
The isomeric excitation energy needed for the first method was previously measured via a two $\gamma$-ray internal transition to the ground-state proceeding through a 630.5 keV (M4) $\rightarrow$ 111.0 keV (M1) cascade[51], and was later confirmed in subsequent experiments which yielded 630.3 keV $\rightarrow$ 111.3 keV [52] and 630.5$\pm$2.6 keV $\rightarrow$ 111.3$\pm$1.1 keV by [53]. The 149mEr isomer is now established to lie 741.8$\pm$0.2 keV above the ground-state according to the NUBASE 2020 evaluation [54].
@@
This excitation energy is consistent with our recent result in which the newly measured mass of 150Yb was used to support a theoretical explanation for the remarkably consistent first excitation energy (approximately 750 keV) in the isotonic chain containing 149Er [19]. This has been attributed to a clustering of single-proton energy levels for the mildly deformed nuclei at $N=81$ [19] resulting from the single hole in the $h_{11/2}$ neutron orbital common to 149Er and many of its isotones [51]. This provides increased confidence in the overall reliability of the established 149mEr excitation energy, from which our mass measurement of 149mEr yields a 149gEr mass excess of -53599$\pm$25 keV, a 143 keV increase from the AME2020 value when using the mass of the long-lived 741.69$\pm$0.23 keV isomer (9.6$\pm$0.6 s) and its presently accepted excitation energy [55] instead of directly measuring the ground-state.
@@
However, the present dataset also enabled a direct measurement of the ground-state mass by utilizing the lower abundance, overlapping TOF peak of the ground-state and employing the hG procedure. Directly fitting the lower abundance 149gEr (as in Figure 1) with the hG function calibrated to the shape of the 149Dy peak yields a mass excess of -53584$\pm$47 keV, an increase in mass of 158$\pm$55 keV from the AME2020 value. The excitation energy of 149mEr can then be determined from the difference between the measured ground-state and isomer to be 726$\pm$53 keV.
@@
These new masses of the 149Er isomer and ground-state lead us to adopt a revised one-proton separation energy of the ground-state which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay. Both the 149,150Tm masses as well as the revised 149Er value were thus needed for an unambiguous identification of the Tm drip-line, since the new Tm masses alone did not exclude the possibility that 150Tm could also be proton-unbound when using the 149Er value from the AME2020. These measurements confirm that 150Tm is indeed proton-bound, agreeing qualitatively with the Bayesian model averaging predictions of [28] that 149Tm is the first proton-unbound thulium isotope. Furthermore, the proton emission Q-value of 149Tm has now been determined and can be used as an input for decay rate calculations.
@@
While 150Tm was shown to be proton-bound, our measurement of 149gTm establishes it as the first proton-unbound isotope in the thulium chain, having a Q-value of +428 keV when calculated using our Tm measurement and the masses of 148Er and 1H from the AME2020. Using $Q=M_{parent}-\Sigma M_{products}$ gives $Q_{p}=M_{{}^{149}Tm}-M_{{}^{148}Er}-M_{{}^{1}\textrm{H}}=428(35)\textrm{keV}$, a decay energy that is on the high side of the AME extrapolation of 250$\pm$201 keV. This value can be used in estimations of the proton emission half-life (e.g. [4, 5]).
@@
Figure 2: The one-proton separation energy ($S_{p}$) of Tm isotopes calculated from the present measurements of masses in the Tm ($Z=69$) and Er ($Z=68$) isotopic chains is shown in blue and connected by solid lines (The 150Tm data point from experiment #1 is offset horizontally for clarity). This includes our new mass of 149Er at $N=81$, but masses from the 2020 AME [55] were used for 1H, 148Er ($N=80$), and 155Er ($N=87$) in calculating $S_{p}$ of Tm. Theory calculations of $S_{p}$ are shown for comparison. Note that the unresolved admixtures of isomeric states of unknown excitation energy in 151,152,154Tm ($N=82,83,85$) may artificially lower their separation energy.
Aside from BSkG3 [26], the models in Figure 2 systematically over-predict $S_{p}$. Thus the present measurements demonstrate that 149Tm is the first proton-unbound nucleus, rather than 148 or 150Tm.
Evolution of the $N=82$ neutron shell
@@
The present work measures the $\Delta_{2N}$ of Tm at the N=82 shell closure and shows that it is, in fact, the first isotone to depend on the binding energy of a proton-unbound nucleus (149Tm). These measurements indicate the persistence of the $N=82$ shell closure, which is consistent with the recent corresponding measurement of Yb [19]. Tm continues the trend in the shell gap of the more stable $N=82$ nuclei to have a more rapid change in binding energy at even-N nuclei.
Figure 3: Comparison of models and experimental data for the empirical shell gap ($\Delta_{2N}$) split into two figures for clarity. The mass data from [62] and [19] was added for $Z=49$ and $Z=90$ respectively. Density Functional Theory and “liquid drop” models are shown.
@@ IV Conclusion and Outlook
The masses of neutron deficient Tm and Er isotopes obtained in this work identify the precise location of the proton drip-line in the Tm isotopic chain. The transition from positive to negative proton separation energy is observed between 149gTm and 150gTm. Our updated mass of 149Er directly impacts the proton-bound nature of 150Tm since it would be the daughter nucleus of the proton emission if this decay mode were possible. The well-understood composition of the mass spectra provides confidence in our identification of isotopes using the MR-TOF-MS. The masses of two previously unmeasured Tm isotopes were measured in this experiment. Additonally, an update to the previously measured mass of 149mEr and a first direct measurement of the ground-state 149gEr determined them to be $\sim$150 keV heavier than the current literature values, thereby establishing 150gTm as proton-bound.
Together, our masses of 149gTm, 150gTm, and 149gEr establish that 149gTm, having 20 fewer neutrons than the only stable isotope of Tm (169Tm, $N=100$), is the first proton-unbound ground-state in this isotopic chain. This agrees with the drip-line prediction obtained from a recent Bayesian analysis of various theoretical models [28].
We directly measured the mass of 149gEr by performing hyperEMG fits on the time-of-flight spectrum obtained with TITAN’s MR-TOF-MS. Higher precision can be achieved by indirectly determining the ground-state mass from the isomer and the previously measured excitation energy (recent observations in the trends in the excitation energies of the (11/2-) isomers [19] have added credibility to this approach). This indirect approach is how the current literature measurement was obtained.
Our measured mass values provide valuable benchmarks for the models used to predict the binding energies of yet-unmeasured nuclei, allowing them to be compared and weighted appropriately.This experimental result can now provide useful feedback to the various theoretical models that use the proton drip-line as a benchmark for their predictive capabilities. As next-generation radioactive ion beam facilities begin to come online, our update to the literature masses highlights the importance of measuring and publishing not only the masses of previously unmeasured isotopes, but also of confirming the mass values in literature which provide integral inputs to models of both physical processes and nuclear properties.
This measurement highlights the challenges that can arise when identifying and measuring masses of heavier exotic species due to the abundance of isomers and possible isobars. Mass values taken from the AME2020 should be used with caution when used for training data or weighting of models.

### v3

@@ III Discussion
The Er isotopic chain was measured alongside Tm during experiment #2 from 154Er to 149Er. During this second measurement, a discrepancy in the mass of 149gEr relative to the literature value was found which significantly affects the one-proton separation energy ($S_{p}$) of the 150Tm ground state. Together, the newly-measured 150Tm mass and the AME2020 mass of 149gEr results in a $S_{p}$ that is well within $1\sigma$ of zero, but using both updated masses yields a positive proton separation energy consistent with proton-bound 150Tm ($2\sigma$).
@@
As with 151Yb, for which the ISAC target was seen to preferentially deliver the spin 11/2 isomer over the ground state, 149Er, having two fewer protons, similarly appeared to favour the excited state. We therefore fit the two-component 149Er peak using two hyperEMG functions having the same parameters as the 149Dy peak; this yields mass values for both the dominant isomeric state ($t_{1/2}=8.9\pm$0.2 s), as well as the ground state ($t_{1/2}=4\pm$2 s) which shows up as an excess within the low-mass tail of the isomer peak.
@@
The ground state mass of 149Er can be determined by two different methods using our data. One approach is to extract it from the measured isomer mass and the known excitation energy as was previously done in an experiment using Schottky mass spectrometry [14], which reported an uncertainty of $\pm$28 keV. The other method is to obtain the mass directly from the fit of the ground state within the tail of the isomer using the known peak shape.
@@
The first approach obtains the mass of 149gEr using the known isomeric excitation energy which was initially measured via a two $\gamma$-ray internal transition to the ground state proceeding through a 630.5 keV (M4) $\rightarrow$ 111.0 keV (M1) cascade [58] This was later confirmed in subsequent experiments which yielded 630.3 keV $\rightarrow$ 111.3 keV [59] and 630.5$\pm$2.6 keV $\rightarrow$ 111.3$\pm$1.1 keV by [60]. The 149mEr isomer is established to lie 741.8$\pm$0.2 keV above the ground state according to the NUBASE 2020 evaluation [51].
@@
This excitation energy is consistent with our recent result in which the newly measured masses of 151g+mYb were used to support a theoretical explanation for the remarkably consistent first excitation energy (approximately 750 keV) in the isotonic chain containing 149Er [15]. This has been attributed to a clustering of single-proton energy levels for the mildly deformed nuclei at $N=81$ [15] resulting from the single hole in the $h_{11/2}$ neutron orbital common to 149Er and many of its isotones [58]. This provides increased confidence in the reliability of the established excitation energy of the long-lived 741.69$\pm$0.23 keV isomer 149mEr (9.6$\pm$0.6 s). Using this value with our current mass measurement of 149mEr to indirectly determine the ground state mass yields a 149gEr mass excess of -53599$\pm$25 keV, a 143 keV increase from AME2020.
@@
However, the present dataset also enables a more direct measurement of the ground state mass by utilizing the lower abundance, overlapping TOF peak of the ground state of 149Er and employing a hyperEMG fitting procedure. Directly fitting the lower abundance 149gEr (as in Figure 1) with the hyperEMG function calibrated to the shape of the 149Dy peak yields a mass excess of -53584$\pm$47 keV, an increase in mass of 158$\pm$55 keV from the AME2020 value. The excitation energy of 149mEr can then be determined from the difference between our measurements of ground state and isomer to be 726$\pm$53 keV.
@@
These new masses of the 149Er isomer and ground state lead us to adopt a revised one-proton separation energy of the ground state which directly impacts the location of the proton drip-line in Tm, since the proton separation energy of a nucleus depends on both the mass of the candidate proton emitter as well as on the mass of the daughter nucleus of the decay. Both the 149,150Tm masses as well as the revised 149Er value were thus needed for an unambiguous identification of the Tm drip-line, since the new Tm masses alone did not exclude the possibility that 150Tm could also be proton-unbound when using the 149Er value from the AME2020. These measurements confirm that 150Tm is indeed proton-bound, agreeing qualitatively with the Bayesian model averaging predictions of [30] that 149Tm is the first proton-unbound thulium isotope. Furthermore, the proton emission Q-value of 149Tm has been determined and can now be used as an input for decay rate calculations.
@@
Figure 4: The one-proton separation energy ($S_{p}$) of Tm isotopes calculated from the present measurements of masses in the Tm ($Z=69$) and Er ($Z=68$) isotopic chains is shown in red and connected by solid lines (The 150Tm data point from experiment #1 is offset horizontally for clarity). This includes our new mass of 149Er at $N=81$, but masses from the 2020 AME [61] were used for 1H, 148Er ($N=80$), and 155Er ($N=87$) in calculating $S_{p}$ for Tm. Theory calculations of $S_{p}$ are shown for comparison. Corrections were made for measurements possibly containing unresolved, long-lived, isomeric states (including the extrapolated isomers in 149,150Tm) but do not impact the conclusion that 150Tm is the last proton-bound Tm isotope.
Aside from BSkG3 [28], the models in Figure 4 systematically over-predict $S_{p}$. Thus the present measurements demonstrate that 149Tm is the first proton-unbound nucleus, rather than 148Tm or 150Tm.
Evolution of the $N=82$ Neutron Shell Closure
@@
The present work measures the $\Delta_{2N}$ of Tm at the N=82 shell closure and shows that it is, in fact, the first isotone for which $\Delta_{2N}$ depends on the binding energy of a proton-unbound nucleus (149Tm). These measurements indicate the persistence of the $N=82$ shell closure, which is consistent with the recent corresponding measurement of Yb [15]. Furthermore, the empirical shell gap at Z=69 continues the trend seen in the more stable $N=82$ nuclei such as Z=63 and 67 of having a local maximum in $\Delta_{2N}$ at the odd-Z nuclei.
Figure 5: Comparison of models and experimental data for the empirical shell gap ($\Delta_{2N}$) split into two figures for clarity. The mass data from [68] and [15] was added for $Z=49$ and $Z=70$ respectively. A selection of Density Functional Theory and “liquid drop” models are shown.
@@ IV Conclusion and Outlook
The masses of neutron deficient Tm and Er isotopes obtained in this work identify the precise location of the proton drip-line in the Tm isotopic chain. The transition from positive to negative proton separation energy is observed between 149gTm and 150gTm. Our updated mass of 149Er directly impacts the proton-bound nature of 150Tm since it would be the daughter nucleus of the proton emission if this decay mode were possible. The masses of two previously unmeasured Tm isotopes were measured in this experiment. Additonally, an update to the previously measured mass of 149mEr and a first direct measurement of the ground state 149gEr determined them to be $\sim$150 keV heavier than the current literature values, thereby establishing 150gTm as proton-bound.
The well-understood composition of the mass spectra provides confidence in our identification of isotopes using the MR-TOF-MS. The mass of 149gEr was determined directly by performing hyperEMG fits on the time-of-flight spectrum obtained with TITAN’s MR-TOF-MS and indirectly through the 149mEr isomer. Together, our masses of 149gTm, 150gTm, and 149gEr establish that 149gTm, having 20 fewer neutrons than the only stable isotope of Tm (169Tm, $N=100$), is the first proton-unbound ground state in this isotopic chain. This agrees with the drip-line prediction obtained from a recent Bayesian analysis of various theoretical models [30].
Our update to the literature masses also highlights the importance of measuring and reporting not only the masses of previously unmeasured isotopes but also confirmations of literature mass values which provide input for various models and predictions. Heavier exotic species pose challenges that arise from the abundance of isomers and isobaric molecules that must be identified and, therefore, must be treated with caution. The mass values presented here provide a new benchmark for the models used to predict the binding energies of unmeasured nuclei in the neutron-deficient region near N=82 and establish the Z=69 proton drip-line.
@@ V Acknowledgements
Table 2: Data from Table 1 without correction for unresolved isomers. Nuclides with a * are known or expected to have a long-lived isomer that could not be resolved in this experiment. No correction is made here and no additional uncertainty added to account for the admixture.
[…]
157g+mLu 157Yb 1.00004820(16) -46373 (25) -46440 (12)
157Tm 157Yb 0.99996379(18) -58714 (28) -58709 (28)
156Tm 156Yb 0.99997554(15) -56818 (24) -56834 (14)
155g+mTm * 155Er 1.00003889(15) -56596 (23) -56626 (10)
155Yb 155Er 1.00008157(15) -50437 (23) -50503 (17)
154g+mTm * 154Dy 1.00011179(18) -54366 (26) -54427 (14)
154Er 154Dy 1.00005458(16) -62569 (24) -62605 (5)
153g+mTm * 153Dy 1.00010676(36) -53935 (52) -53973 (12)
153Er 153Dy 1.00006064(34) -60505 (48) -60467 (9)
152g+mTm * 152Dy 1.00013096(23) -51585 (33) -51720 (50)
152Er 152Dy 1.00006789(18) -60510 (26) -60500 (9)
151g+mTm * 151Dy 1.00012821(17) -50728 (24) -50772 (19)
151Er 151Dy 1.00007472(15) -58248 (22) -58266 (16)
150Tm * 150Dy 1.00016386(23) -46427 (33) -46490 (200#)
150Er 150Dy 1.00008237(16) -57807 (23) -57831 (17)
149Tm * 149Dy 1.00017251(24) -43762 (34) -43940 (200#)
149gEr 149Dy 1.00010171(33) -53584 (47) -53742 (28)
149mEr 149Dy 1.00010695(17) -52857 (25) -53000 (28)
150Tm (Exp 1) * 150Dy 1.00016407(45) -46397 (63) -46490 (200#)